OpenStudy (anonymous):

Will medal for help!! y=4x and y+x=5 A. (-4,1) B. (1,4) C. (-3,2) D. (2,3)

OpenStudy (janu16):

B

OpenStudy (janu16):

Do you want work?

OpenStudy (janu16):

y = 4x Equation(1) x+y=5 Equation (2) Sub. (1) into 2: x + 4x = 5 5x = 5 x = 1 Sub. into (2) x + y = 5 1 + y = 5 y = 4